OpenStudy (anonymous):

A 5.09 mol sample of freon gas was placed in a balloon. Adding 3.50 mol of freon gas to the balloon increased its volume to 23.9 L. What was the initial volume of the balloon?

OpenStudy (joannablackwelder):

Are you familiar with the formula n1/V1=n2/V2?

OpenStudy (joannablackwelder):

For this problem, n1=5.09 mol, V1 is what we are looking for, n2=5.09+3.50 mol, and V2=V1+23.9L.

OpenStudy (anonymous):

Thank you

OpenStudy (joannablackwelder):

No worries. Solving for V1 will take a bit of algebra. Happy studies! :)